# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v1

@@ 3.2 One-dimensional Korteweg-de Vries (KdV) Equation
Figure 5 presents predictions from different frameworks on a representative sample. Both the full rollout and autoregressive DeepONet models fail to capture the dispersive, periodic dynamics of the solution. The full rollout approach, which treats time steps independently, lacks temporal continuity, while the autoregressive model suffers from compounding errors during recursive inference. In contrast, TI-DeepONet and TI(L)-DeepONet closely match the ground truth, even in the extrapolation regime. The adaptive RK4 scheme in TI(L)-DeepONet yields similar accuracy as TI-DeepONet. Figures 3(c) and (d) show the relative $L_{2}$ error over time. The autoregressive DeepONet exhibits rapid error growth beyond $t=2.5$, while the full rollout model shows non-monotonic behavior due to its limited ability to model coherent temporal evolution. In comparison, the TI-based models maintain low and stable errors across the prediction window. At $t=5$, TI(L)-DeepONet and TI-DeepONet achieve final errors of approximately 20% and 19.4%, respectively, 4-5$\times$ lower than baseline methods. The performance difference between TI-DeepONet with AB2/AM3 and TI(L)-DeepONet is marginal, indicating that both are reliable choices for long-term modeling of PDEs similar to the KdV equation.
3.3 Two-dimensional Burgers’ Equation
To assess performance on high-dimensional spatiotemporal dynamics, we consider the 2D Burgers’ equation, defined as:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+u\frac{\partial u}{\partial y}=\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),~{}~{}~{}~{}\forall~{}(x,y,t)\in[0,1]^{2}\times[0,1],$ (7)
where $u(x,y,t)$ is a scalar field and $\nu=0.01$ is the kinematic viscosity. The IC, $u(x,y,0)=s(x,y)$ is sampled from 2D Gaussian random fields, and periodic Dirichlet and Neumann BCs are applied in both $x$ and $y$ directions.
@@ 3.3 Two-dimensional Burgers’ Equation
Figure 6: 2D Burgers’ equation: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated with a representative sample.
@@
Figure 6 presents the spatial error plots in the training and the extrapolation regimes, demonstrating the time integrator models to be significantly more accurate than the baseline models. The solution plot for the same sample is presented in Figure A2. Figure 3 (e) and (f) compares the relative $L_{2}$ error trends over the entire temporal domain. The autoregressive DeepONet rapidly accumulates error, exceeding $85\%$ by the final time step. The full rollout DeepONet initially performs better, but its fixed-basis representation limits its temporal generalization, and errors surpass TI-DeepONet beyond $t=0.5$. Similar to the previous examples, TI(L)-DeepONet maintains slow error growth, with final errors around $15\%$, outperforming all baselines and the TI-DeepONet model in the long-term prediction regime. These results demonstrate the robustness of the time-integrator frameworks in modeling complex spatiotemporal dynamics.
@@ 4 Conclusions
Accurate long-term extrapolation of dynamical systems remains a fundamental challenge for neural operators, particularly due to error accumulation in autoregressive predictions. To address this, we introduce TI-DeepONet and its adaptive variant TI(L)-DeepONet, which integrates classical numerical time integration with an operator learning framework (DeepONet). By reformulating the learning objective to approximate instantaneous time derivatives-subsequently integrated via high-order numerical schemes, our frameworks preserve the causal structure of dynamical systems while enhancing stability and accuracy in long-term forecasting. Furthermore, the operator learning paradigm enables continuous spatiotemporal solutions for parametrized initial conditions, thereby enabling the construction of an efficient and reliable emulator for complex systems. Extensive evaluations across three canonical PDEs (1D/2D Burgers’ and 1D KdV equations) demonstrate the superiority of our approach. TI-DeepONet reduces relative $L_{2}$ extrapolation errors by 81% compared to autoregressive DeepONet and 70% compared to full-rollout methods, while maintaining stable predictions for temporal domains extending up to twice the training interval. The learnable coefficients in TI(L)-DeepONet further refine accuracy by adaptively weighting intermediate integration slopes, enabling solution-specific adjustments that outperform static numerical schemes in all our examples. By embedding temporal causality into the learning process and decoupling training-time integrators (e.g., Runge-Kutta) from inference-time integrators (e.g., Runge-Kutta or Adams-Bashforth), our methodology bridges neural operator flexibility with the robustness of numerical analysis. Among the proposed variants, TI(L)-DeepONet achieves the highest accuracy, albeit with a higher per-iteration computational cost due to its learnable integration coefficients (see SI Section A). However, this increase is offset by its significantly faster convergence to lower training loss, resulting in no net increase in total compute time. This hybrid paradigm offers a reliable pathway for modeling high-dimensional, time-dependent systems, where stability and generalizability are critical.
@@ Acknowledgments
The authors’ research efforts were partly supported by the National Science Foundation (NSF) under Grant No. 2438193 and 2436738. The authors would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the NSF grant number OAC1920103. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding organizations.
@@ Appendix A Limitations: Computational costs
The improved extrapolation accuracy offered by the proposed TI-DeepONet and TI(L)-DeepONet approaches comes at the cost of increased training and inference times. This additional computational overhead stems from the repeated forward and backward passes through the time integrator. However, the convergence of TI-based networks is significantly faster, indicating more efficient optimization of network parameters compared to baseline methods. Table A1 presents a detailed comparison of the training and inference costs for each method across all benchmark problems. Moreover, incorporating adaptive, learnable slope coefficients facilitates convergence to the true solution even faster due to the solution-specific local adaptivity within the time integrator. Training and inference for the 1D Burgers’ and 1D KdV cases were conducted on a single NVIDIA A100 GPU equipped with 40GB memory (Intel Xeon Gold Cascade Lake). The 2D Burgers’ case was performed on a NVIDIA A100 GPU with 80GB memory (Intel Xeon Gold Icy Lake).
@@
Burgers’ (1D) TI(L)-DeepONet (Ours) 256 163.62 2500 5.48
TI-DeepONet AB (Ours) 195.43 5.86
Full rollout 283.94 0.38
Autoregressive 296.09 6.61
KdV (1D) TI(L)-DeepONet (Ours) 256 179.82 1000 4.61
TI-DeepONet AB (Ours) 208.59 4.47
Full rollout 234.98 1.22
Autoregressive 298.44 5.98
Burgers’ (2D) TI(L)-DeepONet (Ours) 128 32.83 1000 7.29
TI-DeepONet AB (Ours) 37.42 6.98
Full rollout 218.68 1.16
Autoregressive 115.72 1.95
@@ Appendix B Statistical Variability
Burgers’ (1D) TI(L)-DeepONet 0.019$\pm$0.003 0.023$\pm$0.003 0.036$\pm$0.004 0.044$\pm$0.005
TI-DeepONet AB 0.031$\pm$0.004 0.037$\pm$0.005 0.057$\pm$0.008 0.070$\pm$0.011
Full Rollout 0.043$\pm$0.002 0.095$\pm$0.004 0.247$\pm$0.028 0.336$\pm$0.053
Autoregressive 0.710$\pm$0.089 1.004$\pm$0.144 1.556$\pm$0.206 1.768$\pm$0.227
KdV (1D) TI(L)-DeepONet 0.054$\pm$0.019 0.065$\pm$0.027 0.075$\pm$0.031 0.111$\pm$0.051
TI-DeepONet AB 0.086$\pm$0.026 0.108$\pm$0.034 0.129$\pm$0.043 0.183$\pm$0.063
Full Rollout 0.776$\pm$0.0004 0.716$\pm$0.0005 0.719$\pm$0.0005 0.795$\pm$0.0007
Autoregressive 0.823$\pm$0.073 0.886$\pm$0.064 0.922$\pm$0.069 0.968$\pm$0.083
Burgers’ (2D) TI(L)-DeepONet 0.111$\pm$0.002 0.121$\pm$0.003 0.143$\pm$0.004 0.155$\pm$0.004
TI-DeepONet AB 0.121$\pm$0.002 0.133$\pm$0.002 0.157$\pm$0.003 0.169$\pm$0.003
Full Rollout 0.131$\pm$0.007 0.194$\pm$0.014 0.357$\pm$0.035 0.453$\pm$0.049
Autoregressive 0.503$\pm$0.017 0.590$\pm$0.024 0.783$\pm$0.052 0.894$\pm$0.075
Across five independent trial runs, TI(L)-DeepONet consistently achieves the lowest extrapolation error across all benchmark problems. In addition to superior accuracy, it also exhibits reduced variance in error estimates, indicating enhanced robustness and reliability. It is noteworthy that in a single run of the 1D KdV problem, TI(L)-DeepONet and TI-DeepONet-AB exhibit comparable performance under a single random initialization (see Table 1). However, the superiority of TI(L)-DeepONet becomes clear when averaged over multiple seeds. As expected, both the full rollout and autoregressive DeepONet models suffer from substantially higher errors, particularly outside the training-time domain. Overall, the TI-based frameworks not only enhance extrapolation accuracy but also offer greater robustness across random initializations, underscoring their effectiveness for modeling time-dependent PDEs.
@@ C.1 One-dimensional Burgers’ Equation
Data preparation:
The training and testing datasets are constructed differently depending on the learning strategy employed. In the full rollout setting, where the goal is to train a DeepONet to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps an initial condition $s(x)$ to the full spatiotemporal solution $u(x,t)$, we consider a set of initial conditions, spatiotemporal query locations, and the corresponding solution field. The data is partitioned into training and testing sets using a train-test split ratio of 0.8. Specifically, we use $N_{\text{train}}=2000$ and $N_{\text{test}}=500$ samples. Additionally, recall that we are interested in assessing the performance of the various DeepONet frameworks in learning the full spatiotemporal field over the entire temporal domain, when it has access to the solution at limited timesteps. In this regard, we consider only half of the temporal domain, i.e., $t_{\text{train}}\in[0,0.5]$ during training. Consequently, the grid for the trunk network is formed by taking a meshgrid over spatiotemporal coordinates, $(x,t)$, where $x\in[0,1]$ and $t=t_{\text{train}}\in[0,0.5]$ during training and $t\in[0,1]$ during testing.
For training the autoregressive DeepONet, TI-DeepONet and TI(L)-DeepONet architectures, we adopt a different strategy for preparing the training and testing datasets. In the case of autoregressive methods, the future states of the system are recursively predicted using the model’s own previous outputs as inputs at each timestep. This approach inherently leads to error accumulation, as prediction errors at one timestep are propagated and potentially amplified in the future timesteps. Understanding and mitigating this issue is a key motivation behind the TI-DeepONet architecture. In this regard, the training dataset is created as follows. The input to the branch network is formed by stacking the solution states from $t=0$ to $t=0.5$, i.e., the dataset $[u^{0}(x),u^{1}(x),\dots,u^{50}(x)]$. Similarly, the output ground truth is formed by stacking the solution states from $t=0.01$ to $t=0.51$, i.e., the dataset $[u^{1}(x),u^{2}(x),\dots,u^{51}(x)]$. The rationale behind arranging the dataset in such a way is that in this case, we aim to model the solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ to learn the mapping: $u^{i}(x)\rightarrow u^{i+1}(x)$, where $u^{i}(x)$ and $u^{i+1}(x)$ are the solution states at the $i^{th}$ and $(i+1)^{th}$ timesteps, respectively. The trunk network processes only the spatial coordinates $x\in[0,1]$.
Training architecture:
Table A3 presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, and batch size. For the TI(L)-DeepONet architecture, we additionally employ an auxiliary feedforward network designed to predict the four RK4 slope coefficients. This network consists of two hidden layers with 32 neurons each, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate scheduling strategy was employed for the auxiliary network predicting the adaptive RK4 slope coefficients. Both train and test losses were monitored at each epoch of training. The final model corresponds to the set of parameters that yielded the lowest test loss over the entire training process. Among all the tested architectures, TI(L)-DeepONet exhibited the fastest convergence, attributed to its capacity to adapt to the local dynamics of the solution. Later in this section, we examine the evolution of the learned RK4 slope coefficients across epochs and compare them to the reference coefficients: $\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}$.
Table A3: 1D Burgers’ equation: Neural network architecture and training settings across different methods.
Method Branch Net Trunk Net Activation Epochs Batch Size
TI(L)-DeepONet [101], [100]$\times$6, [60] [100]$\times$6, [60] Tanh $1\times 10^{5}$ 256
TI-DeepONet [101], [100]$\times$6, [60] [100]$\times$6, [60] Tanh $1\times 10^{5}$ 256
Full rollout [128]$\times$6, [60] [128]$\times$6, [60] Tanh $2\times 10^{5}$ 256
Autoregressive [101], [100]$\times$6, [60] [100]$\times$6, [60] Tanh $1\times 10^{5}$ 256

### v2

@@ 3.2 One-dimensional Korteweg-de Vries (KdV) Equation
Figure 5 presents predictions from different frameworks on a representative sample. Both the full rollout and autoregressive DeepONet models fail to capture the dispersive, periodic dynamics of the solution. The full rollout approach, which treats time steps independently, lacks temporal continuity, while the autoregressive model suffers from compounding errors during recursive inference. Regarding the FNO variants, the full rollout FNO exhibits a noticeable phase shift due to its inability to capture the full spectrum of system frequencies; a consequence of being trained on a reduced temporal domain, thereby incurring significant errors throughout the entire temporal domain. The autoregressive FNO demonstrates particularly interesting behavior: while it achieves near-perfect predictions within the training domain, its performance deteriorates markedly upon entering the extrapolation regime. Here, autoregressive errors accumulate and manifest as spectral instability, characterized by frequency leakage. Specifically, the system’s frequencies become amplified at longer temporal horizons, resulting in erratic oscillations in the solution. Consequently, this approach also fails to reliably capture the inherent dispersive and periodic dynamics of the KdV PDE in the extrapolation domain. In contrast, TI-DeepONet and TI(L)-DeepONet closely match the ground truth, even in the extrapolation regime. The adaptive RK4 scheme in TI(L)-DeepONet yields similar accuracy as TI-DeepONet.
[…]
3.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
The one-dimensional Kuramoto-Sivashinsky (KS) equation is a highly nonlinear PDE that is widely known for its chaotic behavior. It arises in various physical contexts such as thin film dynamics, flame fronts, reaction-diffusion systems, and plasma instabilities. The equation introduces a combination of nonlinear advection, second-order dissipation, and fourth-order dispersion terms, expressed as:
$\dfrac{\partial u}{\partial t}=-u\dfrac{\partial u}{\partial x}-\dfrac{\partial^{2}u}{\partial x^{2}}-\dfrac{\partial^{4}u}{\partial x^{4}},~~~(x,t)\in[0,L]\times(0,\infty)$ (7)
$u(x,t=0)=u_{0}~~~x\in[0,L]$ (8)
Here, $u$ represents the evolving scalar field with an initial profile $u_{0}$. The spatial domain $L$ is equipped with periodic boundary conditions. The nonlinear term $u\dfrac{\partial u}{\partial x}$ accounts for advection, while the dissipative term $\dfrac{\partial^{2}u}{\partial x^{2}}$ contributes to numerical stability, and the high-order dispersive term $\dfrac{\partial^{4}u}{\partial x^{4}}$ models the long-wavelength instability mechanisms that ultimately lead to chaotic dynamics. Further details regarding the data generation process are provided in SI C.3, where we adapt the data generation code from [li2021learning].
@@ 3.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
Figure 6: 1D KS equation: Performance of all the frameworks in the training ($t\in[0,15]$) and the extrapolation regime ($t\in[15,30]$) for a representative sample.
Figure 6 presents the prediction errors incurred by the time-integrator–based neural operators compared to standard baselines, shown for both the training and extrapolation temporal domains for a representative sample. In the training regime, DON FR and FNO AR accurately recover the solution profile, whereas DON AR and FNO FR exhibit relatively higher errors. However, in the extrapolation domain, several interesting behaviors emerge for the standard NO baselines. DON FR, DON AR, and FNO FR incur substantial errors once predictions extend beyond the training horizon, with these errors growing progressively as extrapolation time increases. Notably, FNO AR produces highly noisy and fluctuating solution profiles, indicative of spectral instability or spectral bias artifacts. This behavior manifests as high-frequency leakage, where certain unresolved or weakly controlled high-frequency modes become excessively amplified, resulting in pronounced oscillations in the predicted solution. This phenomenon closely resembles the instability observed in FNO AR predictions for the 1D KdV system. Interestingly, while such behavior was observed only for the KdV (dispersive) and KS (chaotic) systems but not for Burgers (dissipative), we also found that the training of FNO AR was inherently unstable for certain random weight initializations. This suggests that systems exhibiting dynamics beyond standard diffusion may not be well-suited for the FNO autoregressive learning paradigm. In contrast, the TI-based variants demonstrate remarkably higher predictive accuracy in both the training and extrapolation temporal domains. Once again, TI(L)-DeepONet performs slightly better in capturing the unseen chaotic dynamics of the KS equation, which can be attributed to its solution-specific local adaptivity during the learning process.
Figures 3(e) and (f) show the relative $L_{2}$ error growth over time for all frameworks. The autoregressive models, DON AR and FNO AR, are highly susceptible to error compounding under recursive rollout. Consequently, they begin accumulating errors within the initial timesteps, with these errors escalating rapidly in the extrapolation domain to reach values as high as 140% at the final timestep. FNO FR performs marginally better: while it still exhibits monotonically increasing error, the overall magnitude remains slightly lower, though still substantial in absolute terms. DON FR performs well in the training region due to DeepONet’s strong interpolation capabilities; however, its accuracy deteriorates rapidly upon entering the extrapolation domain, where errors grow steeply to reach levels comparable to FNO FR (approximately 80% at the final timestep). In stark contrast, the proposed TI-based variants leverage learned system dynamics by incorporating numerical time integrators within the learning framework. This coupling effectively suppresses error growth in the extrapolation region, even in the presence of chaotic dynamics. Consequently, these models not only achieve high accuracy in the training domain but also maintain relatively stable and controlled error growth throughout the extrapolation interval $t\in[15,30]$.
3.4 Two-dimensional Inviscid Burgers’ Equation
To assess performance on complex high-dimensional spatiotemporal dynamics, we consider the 2D Burgers’ equation, defined as:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+u\frac{\partial u}{\partial y}=\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),\quad\forall~(x,y,t)\in[0,1]^{2}\times[0,1],$ (9)
where $u(x,y,t)$ is a scalar field and $\nu=10^{-4}$ is the kinematic viscosity. The initial condition, $u(x,y,0)=s(x,y)$, is sampled from 2D Gaussian random fields, with periodic boundary conditions applied in both spatial dimensions. Although the equation includes a finite viscosity term, its magnitude is sufficiently small ($\nu=10^{-4}$) as a result of which the system effectively behaves as inviscid, with shock formation expected at longer temporal horizons. This rich dynamical behavior, characterized by the development of sharp discontinuities and complex shock wave interactions, makes this problem an attractive benchmark for evaluating the performance of neural operator surrogates. SI C.4 contains comprehensive details of our data generation approach, which builds upon the code developed by [rosofsky2023applications].
Figure 7: 2D Inviscid Burgers’ equation: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated with a representative sample.
@@ 3.4 Two-dimensional Inviscid Burgers’ Equation
Figure 7 presents the spatial error plots in both training and extrapolation regimes, demonstrating that the time integrator models overall achieve higher accuracy than the baseline models. The solution plot for the same sample is presented in Figure A2. Figures 3(g) and (h) compare the relative $L_{2}$ error trends over the entire temporal domain. Notably, we reduced the training temporal domain to $t\in[0,0.33]$ with the extrapolation domain spanning $t\in[0.33,1.0]$, yielding the most challenging case in this work; all networks must perform $3\times$ extrapolation with no prior knowledge of shock formation at longer temporal horizons (no training data contained shocks for any framework). The autoregressive DeepONet rapidly accumulates error, exceeding 130% by the final time step. The full rollout DeepONet initially performs better, but its fixed-basis representation limits temporal generalization, with errors surpassing TI-DeepONet beyond $t=0.5$. Regarding the FNO variants, FNO FR fails catastrophically in capturing global system frequencies due to insufficient solution information across all timesteps, yielding completely unstable solution fields at longer temporal horizons with errors reaching 226% at the final timestep. FNO AR performs significantly better than DON AR, DON FR, and FNO FR, even outperforming the TI-based variants within the training temporal domain. However, due to its inherent sensitivity to error compounding during autoregressive inference, its excellent training domain accuracy (as low as 8% error) rapidly degrades in the extrapolation regime, reaching over 72% error at the final timestep, surpassing DON FR, which saturates at 52%. Consistent with previous examples, TI(L)-DeepONet maintains relatively slower and stable error growth, achieving final errors around 32% and outperforming all baselines as well as TI-DeepONet in the long-term prediction regime. These results demonstrate the robustness of time-integrator frameworks in modeling complex spatiotemporal stiff dynamics.
4 Choice of Inference Timestep
In our primary evaluations, TI-DeepONet was trained and tested using the same temporal resolution, i.e., the training and inference timesteps ($\Delta t$) were matched. However, the architecture’s compatibility with higher-order numerical integrators suggests that, in principle, inference may be conducted using timesteps different from those used during training. As long as the chosen $\Delta t$ satisfies the Courant–Friedrichs–Lewy (CFL) stability condition, the resulting time integration should remain stable. This flexibility stems from the fact that TI-DeepONet mirrors traditional numerical solvers: both employ explicit time-stepping schemes, with TI-DeepONet approximating the right-hand side of the PDE via a neural operator rather than finite-difference or finite-element discretizations of spatial derivatives.
To investigate this capability, we performed preliminary experiments on the 1D Burgers equation, evaluating TI-DeepONet at three inference timesteps, $\Delta t\in{0.1,0.01,0.001}$, while training was performed at $\Delta t=0.01$. Notably, TI-DeepONet preserved comparable $L_{2}$ error even when evaluated with a significantly coarser timestep ($\Delta t=0.1$) than that used during training. This finding suggests the potential for developing more efficient surrogates that accelerate inference without additional training costs.
@@ 5 Conclusions
Accurate long-term extrapolation of dynamical systems remains a fundamental challenge for neural operators, particularly due to error accumulation in autoregressive predictions. To address this, we introduce TI-DeepONet and its adaptive variant TI(L)-DeepONet, which integrates classical numerical time integration with an operator learning framework (DeepONet). By reformulating the learning objective to approximate instantaneous time derivatives-subsequently integrated via high-order numerical schemes, our frameworks preserve the causal structure of dynamical systems while enhancing stability and accuracy in long-term forecasting. Furthermore, the operator learning paradigm enables continuous spatiotemporal solutions for parametrized initial conditions, thereby enabling the construction of an efficient and reliable emulator for complex systems. Extensive evaluations across four canonical PDEs (1D/2D Burgers’, 1D KdV, and 1D KS equations) demonstrate the superiority of our approach. TI-DeepONet reduces relative $L_{2}$ extrapolation errors by 81% compared to autoregressive DeepONet and 70% compared to full-rollout methods, while maintaining stable predictions for temporal domains extending up to twice the training interval. The learnable coefficients in TI(L)-DeepONet further refine accuracy by adaptively weighting intermediate integration slopes, enabling solution-specific adjustments that outperform static numerical schemes in all our examples. By embedding temporal causality into the learning process and decoupling training-time integrators (e.g., Runge-Kutta) from inference-time integrators (e.g., Runge-Kutta or Adams-Bashforth), our methodology bridges neural operator flexibility with the robustness of numerical analysis. Among the proposed variants, TI(L)-DeepONet achieves the highest accuracy, albeit with a higher per-iteration computational cost due to its learnable integration coefficients (see SI Section A). However, this increase is offset by its significantly faster convergence to lower training loss, resulting in no net increase in total compute time. This hybrid paradigm offers a reliable pathway for modeling high-dimensional, time-dependent systems, where stability and generalizability are critical.
@@ Acknowledgments
The authors’ research efforts were partly supported by the National Science Foundation (NSF) under Grant No. 2438193 and 2436738. The authors would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the NSF grant number OAC1920103. The authors would also like to acknowledge Rajyasri Roy, who investigated the flexibility of the TI-DeepONet architecture and analyzed the inference accuracy when using coarser timesteps at inference compared to the finer timesteps used during training for the 1D Burgers equation. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding organizations.
@@ Appendix A Limitations: Computational costs
The improved extrapolation accuracy offered by the proposed TI-DeepONet and TI(L)-DeepONet approaches comes at the cost of increased training and inference times. This additional computational overhead stems from the repeated forward and backward passes through the time integrator. However, the convergence of TI-based networks is significantly faster, indicating more efficient optimization of network parameters compared to baseline methods. Table A1 presents a detailed comparison of the training and inference costs for each method across all benchmark problems. Moreover, incorporating adaptive, learnable slope coefficients facilitates convergence to the true solution even faster due to the solution-specific local adaptivity within the time integrator. Training and inference for the 1D Burgers’, 1D KdV, and 1D KS cases were conducted on a single NVIDIA A100 GPU equipped with 40GB memory (Intel Xeon Gold Cascade Lake). The 2D Burgers’ case was performed on a NVIDIA A100 GPU with 80GB memory (Intel Xeon Gold Icy Lake).
@@
Burgers’ (1D) TI(L)-DON (Ours) 256 163.62 2500 5.48
TI-DON AB (Ours) 195.43 5.86
DON Full rollout 283.94 0.38
DON Autoregressive 296.09 6.61
FNO Full rollout 24.11 3.36
FNO Autoregressive 285.32 2.12
KdV (1D) TI(L)-DON (Ours) 256 179.82 1000 4.61
TI-DON AB (Ours) 208.59 4.47
DON Full rollout 234.98 1.22
DON Autoregressive 298.44 5.98
FNO Full rollout 10.96 6.31
FNO Autoregressive 306.53 1.49
KS (1D) TI(L)-DON (Ours) 128 117.55 3000 7.67
TI-DON AB (Ours) 122.11 8.12
DON Full rollout 22.39 5.26
DON Autoregressive 232.62 7.15
FNO Full rollout 4.52 7.16
FNO Autoregressive 341.34 5.13
Inviscid Burgers’ (2D) TI(L)-DON (Ours) 32 8.91 1000 8.57
TI-DON AB (Ours) 9.85 4.56
DON Full rollout 64.23 2.29
DON Autoregressive 37.71 8.26
FNO Full rollout 2.67 19.85
FNO Autoregressive 97.96 13.66
@@ Appendix B Statistical Variability
[…]
Across five independent trial runs, TI(L)-DeepONet consistently achieves the lowest extrapolation error across all benchmark problems. In addition to superior accuracy, it also exhibits reduced variance in error estimates, indicating enhanced robustness and reliability. As expected, both the full rollout and autoregressive DeepONet models suffer from substantially higher errors, particularly outside the training temporal domain. Similarly, the FNO variants, both full rollout and autoregressive incur higher errors. FNO FR generally exhibits higher errors throughout the entire temporal domain and yields particularly unstable solutions due to its inability to capture all global frequencies of the system. Notably, this limitation also manifests as higher variance in the predictions. For FNO AR, spectral instability causes error spikes in the extrapolation domain, leading to higher solution variance - consistent with our observations highlighted in Sec. 3. Overall, the TI-based frameworks not only enhance extrapolation accuracy but also offer greater robustness across random initializations, underscoring their effectiveness for modeling time-dependent PDEs.
@@ C.1 One-dimensional Burgers’ Equation
The initial condition $s(x)$ is sampled from a Gaussian random field with spectral density $S(k)=\sigma^{2}(\tau^{2}+(2\pi k)^{2})^{-\gamma}$, where $\sigma=25$, $\tau=5$, and $\gamma=4$. This ensures $s(x)$ is periodic on $x\in[0,1]$. The corresponding kernel is expressed via inverse Fourier transform as $K(\mathbf{x,x^{\prime}})=\int_{-\infty}^{\infty}S(k)e^{2\pi ik(\mathbf{x}-\mathbf{x^{\prime}})}dk$. We discretize the spatiotemporal domain using 101 grid points along each dimension.
Data preparation:
The training and testing datasets are constructed differently depending on the learning strategy employed. In the full rollout setting, the goal is to train a DeepONet/FNO to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps an initial condition $s(x)$ to the full spatiotemporal solution $u(x,t)$. For this purpose, we consider a set of initial conditions, spatiotemporal query locations, and the corresponding solution fields. The data is partitioned into training and testing sets using a train-test split ratio of 0.8, yielding $N_{\text{train}}=2000$ and $N_{\text{test}}=500$ samples. Importantly, we aim to assess the performance of various frameworks in learning the full spatiotemporal field over the entire temporal domain when given access to solutions at limited timesteps. To this end, we restrict training to only half of the temporal domain, i.e., $t_{\text{train}}\in[0,0.5]$. Consequently, the grid for the trunk network is formed by taking a meshgrid over spatiotemporal coordinates $(x,t)$, where $x\in[0,1]$ and $t\in[0,0.5]$ during training, while $t\in[0,1]$ during testing. This same meshgrid is also employed for computing the 2D Fourier Transform in FNO full rollout.
For training the autoregressive DeepONet, autoregressive FNO, TI-DeepONet, and TI(L)-DeepONet architectures, we adopt a different strategy for preparing the training and testing datasets. In autoregressive methods, future states of the system are recursively predicted using the model’s own previous outputs as inputs at each timestep. This approach inherently leads to error accumulation, as prediction errors at one timestep propagate and potentially amplify at subsequent timesteps. Understanding and mitigating this issue is a key motivation behind the TI-DeepONet architecture. To this end, the training dataset is created as follows: the input to the branch network consists of stacked solution states from $t=0$ to $t=0.5$, i.e., $u^{0}(x),u^{1}(x),\ldots,u^{50}(x)]$. Similarly, the output ground truth consists of stacked solution states from $t=0.01$ to $t=0.51$, i.e., $u^{1}(x),u^{2}(x),\ldots,u^{51}(x)]$. This arrangement enables the solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ to learn the mapping $u^{i}(x)\rightarrow u^{i+1}(x)$, where $u^{i}(x)$ and $u^{i+1}(x)$ represent the solution states at the $i^{\text{th}}$ and $(i+1)^{\text{th}}$ timesteps, respectively. The trunk network processes only the spatial coordinates $x\in[0,1]$. For autoregressive FNO, the inputs and outputs are distinct $u^{i}(x)$ and $u^{i+1}(x)$ solution pairs, and the Fourier transform is one-dimensional and is applied only in the $x$-dimension.
Training architecture:
Tables A3 and A4 present the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). For the TI(L)-DeepONet architecture, we additionally employ an auxiliary feedforward network designed to predict the four RK4 slope coefficients. This network consists of two hidden layers with 32 neurons each, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate scheduling strategy was employed for the auxiliary network predicting the adaptive RK4 slope coefficients. Notably for FNO AR, the learning rate schedule also employed exponential decay but by a factor of 0.96 every 2000 steps. Both training and test losses were monitored at each epoch. The final model corresponds to the set of parameters that yielded the lowest test loss over the entire training process. Among all tested architectures, TI(L)-DeepONet exhibited the fastest convergence, attributed to its capacity to adapt to the local dynamics of the solution. Later in this section, we examine the evolution of the learned RK4 slope coefficients across epochs and compare them to the reference coefficients: $\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}$.
Table A3: 1D Burgers’ equation: Neural network architecture and training settings across different DeepONet-based methods.
Method Branch Net Trunk Net Activation Epochs Batch Size
@@
[…]
Table A4: 1D Burgers’ equation: Neural network architecture and training settings across different FNO-based methods.
Method Fourier Blocks Modes Hidden Dim Activation Epochs Batch Size
FNO AR 6 24 64 GELU 30000 128
FNO FR 4 16 32 GELU 5000 64